OpenStudy (anonymous):

If f(4) = 0 and f(6) = 6, which of the following could represent f(x) ?

OpenStudy (anonymous):

y/(x-4) = (6-0)/(6-4)=3 y=3x-12

OpenStudy (anonymous):

thanks. but is there an equation for it?

OpenStudy (anonymous):

yes, y=3x-12 is the equation for it.With different values of x, you get different values of y

OpenStudy (anonymous):

like if x=0,y would be -12. If x=1,y=3*1-12=-9

OpenStudy (anonymous):

i see. i just have to get used to the type of problem. thats all :)